OpenStudy (anonymous):

6x^2=y-1 is symmetric to which axis or line

OpenStudy (anonymous):

the y-axis... the axis of this parabola is the y-axis...

OpenStudy (jlastino):

Because when you change x to -x it would still be equal to the original equation 6(-x)^2=y-1 6x^2=y-1